# Coordinated UAV Beamforming and Control for Directional Jamming and Nulling

### v1

@@ IV-B Receding Horizon Implementation Against a Moving Eavesdropper
Subsequently, we apply the scheme from the previous subsection to a scenario where both the client and the eavesdropper are in motion. We assume the client and eavesdropper are initially located at $6000~0]^{\mathrm{T}}~\mathrm{m}$ and $10000~0]^{\mathrm{T}}~\mathrm{m}$, respectively, and follow northbound trajectories. To account for their changing positions, the UAV replans its trajectory and beamforming every $20$ seconds, with a planning horizon of $t_{f}=150~\textrm{sec}$, and for a total of $1000$ seconds (i.e., 50 replanning instances). Moreover, to ensure that no jamming power reaches the client due to its movement between replanning intervals, the UAV adjusts the beamforming online by recomputing the optimal antenna angle $\theta_{g}(t)$ from the formula (10) and the phase from (6). The rest of the parameters are as previously. Note, each replanning problem took approximately $40~\textrm{ms}$ to solve.
Figure 7 shows the resulting path of the UAV with snapshots of the beampattern, along with the paths of the client and eavesdropper. We notice that the UAV initially moves to the east-southeast, despite the client and eavesdropper moving north. This is because a southeastern trajectory allows the UAV to get better angle separation between the client’s and the eavesdropper’s direction, hence allowing the beampattern $B^{\star}$ on the eavesdropper to increase. Subsequently, the UAV converges to a path between the client and the eavesdropper which maximizes angle separation – and hence $B^{\star}$ on the eavesdropper – while reducing path losses. Finally, the jamming power crosses the $100~\textrm{dBm}$ interference threshold after $62.5$ seconds, and the $90~\textrm{dBm}$ DoS threshold after $280$ seconds.

### v2

@@ IV-B Receding Horizon Implementation Against a Moving Eavesdropper
Subsequently, we apply the scheme from the previous subsection to a scenario where both the client and the eavesdropper are in motion. We assume the client and eavesdropper are initially located at $6000~0]^{\mathrm{T}}~\mathrm{m}$ and $10000~0]^{\mathrm{T}}~\mathrm{m}$, respectively, and follow northbound trajectories. To account for their changing positions, the UAV replans its trajectory and beamforming every $2$ seconds, for a total of $1000$ seconds (i.e., 500 replanning instances). Moreover, to ensure that no jamming power reaches the client due to its movement between replanning intervals, the UAV adjusts the beamforming online by recomputing the optimal antenna angle $\theta_{g}(t)$ from the formula (10) and the phase from (6). The rest of the parameters are as previously. Note, each replanning problem took approximately $15~\textrm{ms}$ to solve.
Figure 7 shows snapshots of the resulting path of the UAV and its beampattern, along with the positions of the client and eavesdropper. We notice that the UAV initially moves to the east-southeast, despite the client and eavesdropper moving north. This is because a southeastern trajectory allows the UAV to get better angle separation between the client’s and the eavesdropper’s direction, hence allowing the beampattern $B^{\star}$ on the eavesdropper to increase. Subsequently, the UAV converges to a path between the client and the eavesdropper which maximizes angle separation – and hence $B^{\star}$ on the eavesdropper – while reducing path losses. Finally, jamming power crosses the $100~\textrm{dBm}$ interference threshold after $42.6$ seconds, and the $90~\textrm{dBm}$ DoS threshold after $88.6$ seconds (with a short drop to $90.4~\textrm{dBm}$ at $t=220~\textrm{sec}$).